# Oxyfluoride glasses obtained through incorporation of CaF2 into photovoltaic cover glass melts

### v1

@@ 4 Conclusion
In this work, we have demonstrated the utilization of end-of-life cover glass from commercial solar panels to produce oxyfluoride glasses by incorporating CaF2. XRF analyses have proven that the final samples are free from contaminants, such as iron, that could introduce color to the material, reducing its value and limiting its potential applications. This confirms that, thanks to the high purity and transparency of the soda-lime cover glass, we could use it to make up to 80% of the total weight of CgCAF glasses. In contrast, conventional flat glass production allows only a small proportion of the material to be recycled. Additionally, the melting temperature of the samples was about 1200oC, significantly lower than the melting temperature of soda-lime glass, which favors a reduced energy consumption and carbon emissions to produce CgCAF samples. The thermal analysis showed a 50-90oC reduction in Tg temperature relative to the soda-lime, the presence of at least two crystallization processes and a pretty good stability of the glass phase, demonstrated by (Tx-Tg) values in the range 100-180oC. XRD confirmed the glassy nature of the samples and indicated a depolymerization process as the CaF2 concentration increases, which is corroborated by mass density and Raman data. In situ XRD measurements in the CgCAF12 sample demonstrated the formation of combeite, diopside, and Ca3SiO5 triclinic phases, besides some unidentified peaks that require further studies. Finally, the UV-VIS-NIR transmittance demonstrated the good optical quality of all samples and the improved UV transmittance window compared to the original cover glass employed in this work. The ensemble of results shows that this family of materials can be explored as glass or be crystallized to produce glass-ceramics that may be tailored for different applications. Further studies should focus on understanding these materials’ crystallization process and incorporating optically active ions. If these oxyfluoride glasses could be used at scale for some application, it could open up a path to reintroduce EOL cover-glass into the economy, improving its value and contributing to the industry’s sustainability.
@@ 5 Acknowledgement
The authors thank CNPq (grants 409475/2021-1, 402473/2023-0 e 304060/2023-2) for the financial support.

### v2

@@ 4 Conclusion
In this work, we have demonstrated the utilization of end-of-life cover glass from commercial solar panels to produce oxyfluoride glasses by incorporating CaF2. XRF analyses have proven that the final samples are free of contaminants, such as iron, that could introduce color to the material, reducing its value and limiting its potential applications. This confirms that, due to the high purity and transparency of the soda-lime cover glass, we could use it to make up to 80% of the total weight of CgCAF glasses. In contrast, conventional flat glass production allows only a small proportion of the material to be recycled. Furthermore, the melting temperature of the samples was about 1200∘C, significantly lower than the melting temperature of soda-lime glass, which favors a reduced energy consumption and carbon emissions to produce CgCAF samples. Thermal analysis showed a 50-90∘C reduction in Tg temperature relative to the soda-lime, the presence of at least two crystallization processes and a pretty good stability of the glass phase, demonstrated by (Tx-Tg) values in the range 100-180∘C. XRPD confirmed the glassy nature of the samples and indicated a depolymerization process as the CaF2 concentration increases, which is corroborated by mass density and Raman data. In situ XRPD measurements in the CgCAF12 sample demonstrated the formation of combeite, diopside, and Ca3SiO5 triclinic phases, besides some unidentified peaks that require further studies. Finally, the UV-VIS-NIR transmittance demonstrated the good optical quality of all samples and the improved UV transmittance window compared to the original cover glass employed in this work. The ensemble of results shows that this family of materials can be explored as glass or be crystallized to produce glass-ceramics that may be tailored for different applications. Further studies should focus on understanding the crystallization process of these materials and incorporating optically active ions. If scaled for practical applications, these oxyfluoride glasses could provide a viable route for the reintegration of end-of-life cover glass into the industrial value chain. Such an approach would not only enhance the economic value of this waste material but also promote circularity and sustainability within the glass and photovoltaic industries.
@@ 5 Acknowledgement
The authors thank CNPq (grants 409475/2021-1, 402473/2023-0 and 304060/2023-2) for the financial support.